Block Island TV currently sells large televisions for $360. It has costs of $290. A competitor is bringing a new large television to market that will sell for $310 Management believes it must lower the price to S310 to compete in the market for large televisions. Marketing believes that the new price w cause sales to increase by 10%, even with a new competitor in the market. Block Island TV sales are currently 110,000 televisions per year What is the target cost if the company wants to maintain its same income level, and marketing is correct (rounded to the nearest cent)? OA. S232.50 O B. $290.00 O C. $246.36 O D. $240.00
Answer #1

Ans: C $246.36

Working:

Current units sold =110,000units

Selling price =$360 and cost per unit =$290

So Current profit=(Sales- costs)*Number of televisions sold

=(360-290)*110,000=$7,700,000

NOW

Selling price per unit = 310$

Units sold now would be =(110,000*110%)=121,000 units

Hence profit required per unit=(7,700,000/121,000)=$63.64(Approx)

Hence Sales - costs=Profit per unit

Hence Target costs=(310-63.64)

=$246.36